If 60 ml of 1F HC2H3O2 is added to 200 ml of 1 F NaC2H3O2, what is the hydronium ion concentration of the resulting solution? Calculate the pH of the solution.


1
Expert's answer
2021-03-29T05:14:57-0400

Molar Mass of HC2H3O2 = 60.052

= 60/60.052 = 1 mol

Molar Mass of NaC2H3O2 = 82.0343

= 200/82.0343 = 2.43 moles

= 1+2.43 = 3.43 m

Hydronium concentration = 1×10^-7 ×3.43

= 3.43 ×10^-7 mol per litre